### v1

# Seeing Through Green: Text-Based Classification and the Firm’s Returns from Green Patents

Seeing Through Green: Text-Based Classification and the Firm’s Returns from Green Patents
Lapo Santarlasci Mail to: lapo.santarlasci@imtlucca.it. Laboratory for the Analysis of Complex Economic Systems (AXES), IMT School for Advanced Studies, Lucca    Armando Rungi Mail to: armando.rungi@imtlucca.it. Laboratory for the Analysis of Complex Economic Systems (AXES), IMT School for Advanced Studies, Lucca    Antonio Zinilli Mail to: antonio.zinilli@cnr.it. National Research Council of Italy (CNR), Research Institute on Sustainable Economic Growth, Unit of Rome
This paper introduces Natural Language Processing for identifying ’true’ green patents from official supporting documents. We start our training on about 12.4 million patents that had been classified as green from previous literature. Thus, we train a simple neural network to enlarge a baseline dictionary through vector representations of expressions related to environmental technologies. After testing, we find that ’true’ green patents represent about 20% of the total of patents classified as green from previous literature. We show heterogeneity by technological classes, and then check that ’true’ green patents are about 1% less cited by following inventions. In the second part of the paper, we test the relationship between patenting and a dashboard of firm-level financial accounts in the European Union. After controlling for reverse causality, we show that holding at least one ’true’ green patent raises sales, market shares, and productivity. If we restrict the analysis to high-novelty ’true’ green patents, we find that they also yield higher profits. Our findings underscore the importance of using text analyses to gauge finer-grained patent classifications that are useful for policymaking in different domains.
Keywords: green patents, text analysis, natural language processing, patent novelty, environmental innovation, firm level
JEL Codes: O31, F64, O34, O35, Q55
1 Introduction
Climate change is no longer a distant externality; it is a binding economic constraint that is already reshaping product markets, capital allocation and regulatory priorities worldwide. In the European Union, for example, the Green Deal, the Fit-for-55 package and a growing array of national carbon-pricing schemes collectively tighten the link between a firm’s environmental footprint and its license to operate. Against this backdrop, green innovation has emerged as a central lever for reconciling competitiveness with decarbonisation goals. Yet a decade of empirical work has produced mixed verdicts on whether green innovation translates into superior firm performance. While several studies document positive stock-market reactions or profitability gains, others find negligible or even negative effects once R&D crowd-out and compliance costs or demand uncertainty are considered (Barbieri et al., 2017; Przychodzen and Przychodzen, 2015).
In this contribution, we start by challenging standard classifications of green patents, i.e., widely used lists such as OECD’s ENV-TECH, the EPO’s Y02/Y04S and WIPO’s IPC Green Inventory assign green status to entire patent subclasses. We believe that previous approaches guarantee coverage but risk conflating genuinely environmentally efficient inventions with conventional technologies that merely happen to sit in an environmentally relevant class. In this paper, we propose the adoption of a Natural Language Processing method based on a simple neural network algorithm that allows us to end up with a dictionary for environmental technologies. Through a so-called continuous bag of words, the algorithm uses a shallow neural network with a single hidden layer. The model learn semantic relationships from patents titles and abstracts, and leverage them to enlarge the seed dictionary derived from large language models, with synonyms. The model computes the 15 most semantically similar expressions for each seed. For our purpose, we train our algorithm on a unique database of about 12,400,000 patents. The result is a selection of what we consider true green patents, which is substantially narrower, i.e., 20%, than the one based on existing classifications.
We believe our proposal is particularly useful in combination with another text-based indicator proposed by Arts et al. (2021), which catches the novelty of patents based on previously unseen keywords and phrases. What we want to avoid with this implementation is for greenwashing to hide in plain sight. As regulatory and investor attention intensifies, firms may strategically patent marginal design tweaks to signal sustainability credentials — a form of greenwashing that inflates the counts of green patents without a corresponding technological progress in environmental protection. Once we consider also the novelty degree of true green patents, we find that they constitute about 49% of the total.
In the second part of our paper, we link patents with European firms, when the latter are owners of intellectual property, to investigate whether there is an association with firms’ competitiveness. Several researchers have already contributed significantly to the study of patents and firm performance. Key contributions include the work of Griliches (1998), who provided early surveys on patent statistics as economic indicators. Other foundational research linked firm patenting activity to stock market valuations, as in Pakes (1985). A notable contribution providing comprehensive evidence on what happens when firms patent, specifically linking patenting to real economic effects like firm growth through new product introductions, is the work by Balasubramanian and Sivadasan (2011). The authors clearly explain how difficult it is to purge different sources of endogeneity in the relationship between patents and firm outcomes. Thus, drawing causal conclusions is a challenge. The decision for a firm to patent is influenced by underlying factors such as the innovation ability, the potential market demand, or strategic market interactions. Consequently, observed positive correlations between patenting activity and firms’ financial accounts are not informative of the direction of causality. See also Exadaktylos et al. (2024), who propose an instrumental variable strategy based on exogenous variation at the level of the patent office. The benefits from legal protection afforded by patents can be confounded with the benefits deriving from the innovative content of the patents.
A final solution to the endogenous linkage between firms and financial accounts is beyond the scope of the present paper. In the next paragraphs, we design a basic identification strategy to show the value of our exercise on true green patents, and compare with broader classifications in the literature. Beyond correlations, we propose to address reverse causality, i.e., when bigger and more productive firms have a higher propensity to generate an innovation and register a patent. Starting from a potential outcomes framework á la Rubin (2005), we implement a propensity score matching that controls for firm size, age, and industry affiliations. Our units are considered treated when they register green patents. Therefore, we approximate a counterfactual scenario of green innovation on economic outcomes, assuming selection on observables. The results are discussed by considering the different economic channels that explain observed positive effects.
We find that holding at least one green patent is associated with more than double sales, a market share that is higher by 38%, and productivity gains by 40%. If we further check for the novelty degree of the green patents, after we interact with the indicator by Arts et al. (2021), we find that high-novelty green patents have almost double the gains on sales. Previous results can be due to a combination of protection of property rights and the innovative content of the invention, both of which provide an explanation for a positive association with firms’ financial accounts (Exadaktylos et al., 2024).
The remainder of the paper is structured as follow: in Section 2 we present the related literature highlighting our contribution and the reason behind our research questions; in Section 3 we present the dataset we constructed and we show some descriptive evidence; in Section 4 we the present our custom NLP classification procedures and its results; in Section 5 we present our empirical strategy and final results, and, finally, in Section 6 some robustness and sensitivity checks.
2 Literature review
Our paper contributes to four different strands of literature. First, it primarily contributes to questioning the effectiveness of broad classification approaches to investigate the environmental scope of patents, e.g., Wagner (2007); Johnstone et al. (2010); Rave et al. (2011); Ghisetti et al. (2021); Block et al. (2025); Rainville et al. (2025). Second, it contributes to the line of studies investigating the association between intellectual property rights and firms’ performance, e.g., Bloom and Van Reenen (2002); Hall et al. (2005); Huang and Li (2017); Amore and Bennedsen (2016); Barbieri et al. (2017); Cai and Li (2018); de Azevedo Rezende et al. (2019); El-Kassar and Singh (2019); Fernandes et al. (2021); Exadaktylos et al. (2024); Portillo-Tarragona et al. (2024); Chirico et al. (2025). Third, we argue that our work complements previous efforts to explore metrics that define the degree of novelty of a patent (Artz et al., 2010; Hirshleifer et al., 2013; Kogan et al., 2017; Chirico et al., 2025). Finally, by understanding when a patent actually contributes to the green transition, we can comment on the limits and consequences of greenwashing, e.g., Kammerer (2009); Berrone et al. (2017), when companies deceptively use intellectual property rights to persuade the public that they are environmental friendly.
According to the World Intellectual Property Organization (WIPO), an invention is “a product or a process that provides a new way of doing something or offers a new technical solution to a problem that surpasses trivial solutions”; while a patent grants an inventor exclusive rights and offers legal protection for such an invention. Generally, a green patent is understood as a granted patent for an invention aimed at reducing negative environmental impacts or generating environmental benefits. The overarching purpose of green innovation — and, accordingly, green patents — is to address environmental goals such as waste reduction, eco-efficiency, and the abatement of emissions (Chirico et al., 2025).
Two primary approaches emerge in the literature for identifying green innovations from patents: the use of patent classification systems and keyword searches within the patent text. The former method relies on established expert systems such as the International Patent Classification (IPC) of WIPO and the Cooperative Patent Classification (CPC) employed by the European Patent Office (EPO) and the United States Patent and Trademark Office (USPTO). Agencies like the OECD, WIPO, and EPO provide dedicated lists of IPC or CPC codes corresponding to green technologies. A patent is deemed green if it falls under these codes — examples include WIPO’s Green Inventory, the EPO’s Y02 list for climate change mitigation technologies, or the OECD’s ENV-TECH classification, which covers a wider set of environmental innovations. Increasingly, IPC and CPC codes have gained popularity for identifying green patents because they are widely available and informed by expert assessments (Block et al., 2025). Many studies therefore classify a patent as green if its codes align with recognized green technology lists, for instance, Barbieri et al. (2020) used bibliographic patents data to locate environment-related patents based on the OECD ENV-TECH classification, while Fabrizi et al. (2018) employed the “Y02” scheme to target climate change mitigation technologies. Among others, Bellucci et al. (2023) relied on a methodology developed by the European Commission’s Joint Research Centre (JRC) to construct global indicators of innovation in clean energy technologies, also using patent classifications. León et al. (2017) focused on patent families and international patent applications through WIPO’s PCT in green energy technologies, referencing WIPO’s IPC Green Inventory. Nonetheless, one acknowledged drawback of approaches based on classification codes is that categories can be overly broad, and to address this, some scholars adopt more granular classifications by considering subclassification codes (Amore and Bennedsen, 2016). Moreover, classification codes can be prone to subjective interpretation by patent applicants and examiners, potentially skewing results if green and non-green technologies are classified differently (Barbieri et al., 2020). Discrepancies among classification systems imply a patent qualifies as green under one system but not under another, and such a lack of harmonisation in identifying green patents within IPC and CPC codes is well documented (Rainville et al., 2025).
A second approach involves mining the text of patents — particularly titles, abstracts, and claims — using environment-related or ecological keywords. Wagner (2007), for example, identified green innovations by scanning patent abstracts for ecological terms, while Ghisetti et al. (2021) constructed a dictionary-based keyword approach informed by the ENV-TECH codes descriptions list to pinpoint green designs and trademarks. The European Union Intellectual Property Office (EUIPO) similarly deployed a method that classifies over 85,000 descriptive terms (within the Nice [^1] The Nice Classification (NCL) is an international classification of goods and services applied for the registration of marks. classification for goods and services) as green or non-green, as determined by expert evaluators (European Union Intellectual Property Office, 2021; Block et al., 2025). Although text mining can capture green innovations that might be overlooked by classification codes, its effectiveness depends on carefully chosen keywords; patents that either do not use these specific terms or use them ambiguously may be incorrectly included or excluded. For example, Nameroff et al. (2004) notably found that “green chemistry” and “green chemical” did not appear in any patent titles, abstracts, or claims at the time of their study, underscoring a key limitation when relying solely on narrow terminology.
Against the previous background, we propose an identification of an environmental technological dictionary after a large-scale training on the widest set of patents that have ever been classified as green by one of the existing classifications. Starting from seed expressions proposed by Large Language Models, we use a Word2vec algorithm (Mikolov, 2013; Mikolov et al., 2013) to identify the combinations of words that will be included in the baseline dictionary. Eventually, we will obtain the subset of the patents that we can consider as true green.
Various tailored strategies have also been proposed. Amore and Bennedsen (2016), for instance, used a detailed classification strategy that combined primary 3-digit USPTO codes with sub-classifications to examine energy patents. Ghisetti et al. (2021) combined IP expertise with a semi-automated procedure to refine an initial set of keywords, thus ensuring relevance for design and product-oriented innovation under the OECD Environmental Technology classification. Other researchers have proposed merging classification-based and keyword-based strategies to increase accuracy (e.g., (Wagner, 2007; Block et al., 2025). Rainville et al. (2025) addressed difficulties in categorising patents tied to the circular economy (CE), demonstrating that existing green patent classifications do not fully capture CE principles. They thus suggested a keyword analysis of patent abstracts to more accurately identify CE-related patents within the CPC scheme, a finding echoed by Portillo-Tarragona et al. (2024), who extended the notion of circular patents beyond waste-related or green patents to those supporting a circular or sharing economy.
Overall, while classification-based methods remain prevalent for pinpointing green patents, thanks to their accessibility and grounding in expert knowledge, debate persists regarding their limitations, since the method chosen can significantly impact both scholarly insights and policy-making recommendations, underlining the importance of methodological rigour and transparency in identifying green patents. Our contribution relates then to the branch of tailored custom strategies, proposing a novel NLP-based classification method, aiming to avoid misclassification problems and to further filter for highly innovative content. This approach’s main goal is to see through the noise of falsely classified green patents and green-washing practices, in order to have a more precise and granular perspective of the green innovation panorama, helping policymakers evaluate economic impacts with greater precision.
Patents are widely viewed as a key output of innovation and can significantly contribute to economic value creation (Artz et al., 2010; Howell, 2017). Nonetheless, patent quality and impact vary considerably; while most patents represent incremental advancements, only a small subset introduces radical innovation (Artz et al., 2010), and empirical evidence points to a positive correlation between higher-quality patents and stronger financial or economic performance (Balasubramanian and Sivadasan, 2011; Hirshleifer et al., 2013; Kogan et al., 2017). To assess this quality, researchers often use patent citations as an indicator of both technological and economic value (Trajtenberg, 1990; Hall et al., 2005; Kogan et al., 2017; Andriosopoulos et al., 2022). Additional measures include patent scope (Barbieri et al., 2017), patent family size (Lanjouw and Mody, 1996; Hall and Helmers, 2013), and the count of non-patent references (Callaert et al., 2006; Hall and Helmers, 2013). Of particular interest is the work by (Arts et al., 2021), which introduces NLP-based measures that often outperform traditional classification and citation-based approaches. These new measures, for example, more accurately identify patents associated with prestigious awards (indicative of high impact) and patents granted by the USPTO but rejected by the EPO and JPO (suggestive of low novelty). Specifically, we believe that it is important to correctly measure green patents and their novelty to avoid greenwashing [^2] Greenwashing refers to “misleading the public to believe that a company or entity is doing more to protect the environment than it is” (United Nations, 2022) practices. Firms may engage in greenwashing to bolster their environmental legitimacy and gain support from consumers and green investors (Berrone et al., 2010, 2017), or they might seek to cultivate a green image (Cai and Li, 2018). While green patents can signal environmental progress to investors (Amore and Bennedsen, 2016), their market valuation depends on firm characteristics and on the specific attributes of the patent (Chirico et al., 2025).
3 Data and descriptive evidence
We source and combine data from PATSTAT, on the one hand, for patent information and from ORBIS, on the other hand, for firm-level data. PATSTAT is a data provider for patent intelligence and statistics; it contains bibliographical and legal events on a global scale. We examine patent families, as they encompass all patent documents linked to a single invention or innovation by the same inventors or assignees. This approach prevents an inflation of patent counts caused by applications for grants in multiple technological classes or jurisdictions. After we work on patents alone, to implement our text analysis, a more limited set of patents is matched with firm-level information from the ORBIS database, which provides comprehensive information on companies around the world. ORBIS is known for its extensive coverage of firm characteristics, including financial accounts, industry classifications, ownership structures, and management data. Given its appealing match of firms and patents, it has been used in previous contributions, e.g., Szücs (2018); Clò et al. (2020); Cappelli et al. (2023); Exadaktylos et al. (2024).
3.1 Patent data
The key variables extracted from PATSTAT include the abstract and the title, providing a text where we can find a summary of the invention used to classify and identify green patents. Additionally, we keep classification codes, specifically the International Patent Classification (IPC) and the Cooperative Patent Classification (CPC), which are used in previous literature to understand green patents by technology classes. Importantly, we can include in our data the filing, publication and granting dates. In particular, these are dates that provide a timeline for the innovation process. These are the events that represent, respectively, the submission of a patent application to the patent office, the publication of such application making it available to the public, and the acceptance by the patent office. Specifically, we acknowledge the filing date as the closest proxy date to the actual date of invention.
To initiate the classification process, we employ three broadly acknowledged algorithms to source patents from PATSTAT: the ENV-TECH, developed by the Organisation for Economic Co-operation and Development (OECD, Haščič and Migotto (2015)), the EPO “Y02/Y04S tagging scheme” (Angelucci et al., 2018) and the WIPO and UNFCCC “IPC Green Inventory” [^3] https://www.wipo.int/classifications/ipc/green-inventory/home. They operate based on the International Patent Classification (IPC) or the Cooperative Patent Classification (CPC), which consider patents as attached to specific technological fields. Table 1 shows an example of IPC codes that the ENV-TECH algorithm links to green patents.
Table 1: Sample of IPC classes used by the ENV-TECH algorithm
IPC Description
B01D Chemical or biological purification of waste gases
B01K Treatment of wastewater
B01N Recycling and recovery of waste materials
B01S Air pollution abatement
  • Note: The table reports examples of IPC classes used by the ENV-TECH algorithm for the classification of patents as “green”.
Although developed independently, the three methodologies are complementary since they cover non perfectly overlapping technological domains, so we use them in combination to obtain a wider coverage for our exercise. Table 2 summarizes the main differences between these methodologies.
Table 2: Classification codes
Developer Methodology Classification codes base
OECD ENV-TECH (OECD) IPC and CPC codes
EPO, UNEP, ICTSD Y02/Y04S (EPO) CPC codes only
WIPO & UNFCCC IPC Green Inventory (WIPO/UNFCCC) IPC codes only
  • Note: The table reports a summary of classifications for the identification of green patents.
Finally, we follow the methodology of Arts et al. (2021) to quantify the novelty and technological significance of each patent based on the lexical contents of patent texts. A patent is relatively more novel if word combinations appear for the first time in abstracts and titles. To do so, we need to consider the historical collection of all granted applications available in PATSTAT. For further details about the methodology, please refer to Appendix A.2. Assessing patent novelty.
3.2 Firm-level data
In the second part of our paper, we reduce the geographical scope of our analysis and study how green patenting is associated with firms’ competitiveness in the European Union. We select all companies with available data to investigate an entire dashboard of firms’ financial accounts in the period 2010-2022. At the beginning, we consider all firms, whether they have patents or not. We end up with a sample of 2,676,515 European firms. Appendix Table B1 shows the number of companies by country, while Appendix Table B2 shows the classification of firms by number of employees. Notably, in Appendix Tables B3, we show that our firm-level match covers all 2-digit industries. When we select only those that have registered patents in our period of analysis, we have a subsample of 15,462 companies. Importantly, among them, only 3,700 have one green patent according to previous classification in the literature, and if we look at the set of firms that have at least one true green patent in the same period, we end up with only 796 firms. For our purpose, we design a dashboard of firms’ outcomes that allows us to understand the gains from green patenting. For each firm, when available, we select sales, market share, labor productivity, capital intensity, ROCE (Return on Capital Employed), EBIT (Earnings Before Interest and Taxes), and TFP (Total Factor Productivity).
Market shares are calculated as the share of sales in a firm’s country and 2-digit NACE Rev. 2 industry. Labor productivity is the ratio between value added and employees, while capital intensity is the ratio between fixed assets and employees. Two measures of profitability, like ROCE and EBIT, enable us to better discuss when profitability gains are statistically significant and why. Finally, TFP can be estimated on a subsample of firms in the manufacturing industries, and we employ the methodology proposed by Ackerberg et al. (2015), which controls for the simultaneity bias between productivity shocks and the choice of production factors.
4 A Natural Language Processing approach to classify true green patents
Although classification codes are widely used for separating green patents, they can lead to cases of misclassification. We report anecdotal evidence in Table 3. The code F01N7 “Exhaust or silencing apparatus, or parts thereof” from the ENV-TECH list can surely be associated with a technical field in which innovation often leads to less polluting engines, as a result of the use of more advanced exhaust systems. But the same code can include an invention for a component that increases engine performance by burning a larger amount of fuel, regardless of environmental concerns.
Table 3: Example IPC classes used by the ENV-TECH algorithm
IPC Description
F01N7 Exhaust or silencing apparatus, or parts thereof
F02M39-71 Fuel-injection apparatus
  • Note: The table reports examples of controversial IPC classes used by the ENV-TECH algorithm for the classification of green patents.
Motivated by previous evidence and many other cases of misclassification, we propose a strategy for the identification of true green patents based on the Natural Language Processing of abstracts and titles. We construct a dictionary for green technologies by means of a basic neural network model, following the Word2Vec algorithm (Mikolov, 2013; Mikolov et al., 2013), which captures semantic relationships between words. We start by extracting from PATSTAT the sub-sample of patents that are classified as green by any of the classification codes proposed in previous literature. Then, we identify a basic set of seed expressions that serve as the foundation for expanding a baseline vocabulary for green technologies. For this first stage, we use ChatGPT and Gemini, two freely available transformer-based chatbots from OpenAI and Google, to produce 70 expressions [^4] See Appendix A.1. Green-patents vocabulary Table A1 for the complete list of seed expressions.. We then merge the two lists and drop duplicates. Thus, we can proceed with the vocabulary expansion by means of Word2Vec [^5] Please refer to Appendix A.3. Word2Vec specification for details about hyperparameter selection embeddings.
In Figure 1, we represent the architecture of the bags-of-words variant of Word2vec that we use. The objective of the algorithm is to maximise the likelihood of predicting a target word given the context words. Thus, the model computes the probability distribution over the entire vocabulary and updates the weights to minimise the prediction error between the actual target word and the predicted words. The output scores are correlated with the likelihood of each word in the vocabulary being the target word given the context. Notably, the bag-of-words variant of the algorithm is trained using backpropagation and stochastic gradient descent to iteratively adjust the weights. Once trained, the weights between the input and the hidden layer (or the hidden layer itself) are used as word embeddings, i.e., they capture semantic relationships between words based on their co-occurrence in the training data.
Figure 1: Neural network architecture
[…]
  • Note: The figure is a representation of the basic neural network for the continuous-bags-of-words (CBOW) variant of the Word2vec algorithm.
In notation let $V$ be the vocabulary size and $d$ be the embedding dimension, which we set at $450$. We define: $W_{in}$$\in$} ${\mathds{R}}^{V\times d}$ as the input weight matrix and $W_{out}$$\in$} ${\mathds{R}}^{d\times V}$ as the output weight matrix. For a target word $w_{t}$ at position $t$ with context window size $C$ we define context words $\{{w_{t-C},\ldots,w_{t-1},w_{t+1},\ldots,w_{t+C}}\}$. As from our hyperparameters’ fine-tuning described in Appendix A.3. Word2Vec specification, we set $C=2$. For each context word ${w_{t-C},\ldots,w_{t-1},w_{t+1},\ldots,w_{t+C}}$, we get their vectors:
$\mathbf{x}_{t+j}\in{0,1}^{V}\quad\text{for }j\in{-C,\ldots,-1,1,\ldots,C}$ (1)
Thus, we compute the average of word embeddings as:
$\mathbf{h}$ $=\frac{1}{2C}\sum_{j=-C,j\neq 0}^{C}\mathbf{W}_{in}\mathbf{x}_{t+j}=\frac{1}{2C}\sum_{j=-C,j\neq 0}^{C}\mathbf{v}_{w_{t+j}}$ (2)
where $\mathbf{v}_{w_{t+j}}$ is the embedding vector for word $w_{t+j}$.
Therefore, we can use $\mathbf{h}$ to predict the target word $w_{t}$ looking at the largest value of the output vector:
$\mathbf{x_{t}}=\mathbf{W_{out}}\mathbf{h}$ (3)
The training process consists, then, of minimising the prediction error of the contextual word forecast. Such a process ends up tuning word embeddings in a way that they represent semantic relationships, given the context. After training Word2Vec, each word $w\in V$ has a learned embedding vector $\mathbf{v}_{w}$ that we can leverage to produce synonyms for each seed expression of our baseline vocabulary. Specifically, we extract the top 15 most semantically similar words [^6] Please note that, at this stage, a manual check might be necessary to exclude a few ambiguous expressions. by means of cosine similarity. In notation, for two words $w_{1}$ and $w_{2}$ embedded in vectors $\mathbf{v}_{w_{1}}$ and $\mathbf{v}_{w_{2}}$, respectively, cosine similarity is:
$\text{sim}(w_{1},w_{2})=\frac{\mathbf{v}_{w_{1}}\cdot\mathbf{v}_{w_{2}}}{|\mathbf{v}_{w_{1}}|\cdot|\mathbf{v}_{w_{2}}|}=\frac{\mathbf{v}_{w_{1}}^{T}\mathbf{v}_{w_{2}}}{\|\mathbf{v}_{w_{1}}\|_{2}\|\mathbf{v}_{w_{2}}\|_{2}}$ (4)
where $\|\mathbf{v}\|_{2}=\sqrt{\sum_{i=1}^{d}v_{i}^{2}}$ is the Euclidean norm.
Our custom vocabulary for environmental technologies is finally used to perform regex matches on patent descriptions, on the sub-sample of patents classified as green by the code-based algorithms, enabling us to identify patents that are more closely aligned with the concept of green innovation. We apply it to abstracts and titles of patents in English [^7] A few patents are not written in English, and we translate them with the help of the Python translator, which systematically leverages Google Translate, after translation, we also adopt NLP techniques to: remove stop words, and apply part-of-speech tagging to retain only nouns, adjectives, verbs, adverbs and proper nouns. We convert to lowercase, lemmatize every word, replace numbers and measures (e.g., LxW or LxWxH) with a generic tag, and remove special characters. Finally, we concatenate titles and abstracts into unique texts
We refer to the matched patent applications as true green patents, flagging out those that unambiguously meet our criteria for environmental relevance. To illustrate the effectiveness of our approach, consider the following extract from a patent abstract that was successfully matched using our custom vocabulary: “The method for producing green fuel comprises a fermentation of vegetable materials, a separation… an anaerobic digestion of said organic material suspension… producing electricity and heat by combusting this biogas, and… supplying energy to at least one step of the method…” This part of the abstract had already been identified as green on the basis of official classifications in existing literature. Our textual matching confirms its environmental relevance by detecting key expressions such as ’green fuel’, ’anaerobic digestion’, ’biogas’ and the reuse of electricity and heat within the process.
Conversely, we also identify cases where the classification codes flagged patents as green, but our textual analysis does not confirm it. For example, patents mentioning generic terms like “carbon” without contextual relevance are discarded unless accompanied by domain-specific expressions (e.g., “carbon capture and storage” or “carbon-based materials” in unrelated industrial applications). These false positives are filtered out by enforcing co-occurrence rules and semantic proximity constraints during the vocabulary refinement and regex matching steps. As an illustration of a non-green case, consider the following extract: “The invention relates to a method for the preparation of metal or metal oxide catalysts that are supported on porous materials. The inventive method is characterised in that it comprises the following steps consisting in: impregnating activated carbon with a catalytically-active phase or with precursors of the catalytically-active phase, shaping a paste, forming structures such as honeycomb or spheres, and subjecting the product to heat treatment to eliminate the activated carbon.”. This abstract has terms such as ’catalyst’ and ’honeycomb structure’, but a closer manual inspection reveals that it merely describes a general process for catalyst preparation. It describes a technical process for the preparation of catalysts without explicitly indicating any environmental application or purpose. We conclude that our procedure makes it possible to identify environmentally relevant patents with greater precision.
4.1 Descriptive evidence
4.1.1 Share of true greens in technological classes
The final result of our NLP approach is a vocabulary tailored to the domain of green technologies. In Appendix Tables A1 A2, A3 and A4, we report a comparison between the number of green patents obtained from previous classifications in each CPC code and the number of true green patents that we obtain with our approach. In Figure 2, we summarize the shares of true greens for each CPC code and their density. On average, we observe a 20% share of the total, but with a dispersion that runs from 0 to 100% [^8] It is important to highlight that a single invention effort can be acknowledged as a patent in more than one technological class, hence descriptive statistics here include multiple counting..
When shifting to a more granular level of classification, we notice how, specifically, CPCs related to “Technologies or applications for mitigation or adaptation against climate change” (Y02), “Electric elements” (H01) and “Physical or chemical processes or apparatus in general” (B01) are the ones with the highest number of green patents if we believe in previous classifications. When we use our approach, CPC Y02 remains the category associated with the largest number of true greens, but then we find “Generation, conversion, or distribution of electric power” (H02) and “Treatment of water, waste water, sewage, or sludge” (C02) in the second and third positions. If we look instead at the share of true green patents we see “Sugar industry” (C13), “Doors, windows, shutters, or roller blinds, in general; ladders” (E06) and “Heating; ranges; ventilating” (F24) in the top 3 rank with shares between 40% and 50%.
Figure 2: Density plot of the shares of true green patents by CPC category.
  • Note: The figure represents the density of the shares of true green patents vs the total of green patents identified by previous classifications.
4.1.2 Technological classes and their specialisation in green patents
The exercise we propose in this section serves a double purpose. First, we want to check which technological classes specialise more in true green patents. Second, we want to check whether our exercise identifies a different pattern of specialisation if compared with traditional classifications. To capture specialisation, we compute the so-called Balassa (1965) index, as follows:
$RCA_{c}=\frac{\frac{G_{c}}{N_{c}}}{\frac{\sum_{c^{\prime}\neq c}G_{c^{\prime}}}{\sum_{c^{\prime}\neq c}N_{c^{\prime}}}},$ (5)
where $G_{c}$ is the number of true green patents issued with in the technological class $c$, $N_{c}$ is the number of non-true-green patents issued in a technological class $c$, $G_{c^{\prime}}$ is the number of true green patents issued in all other classification classes, $N_{c^{\prime}}$ is the number of non-true-green patents issued in all other classes. We believe the RCA index is a powerful measure to catch how much a technological domain is specialised in environmental inventions, also controlling for global patenting trends. In this context, an index value greater than 1 indicates that a technological class is relatively more specialised in green patenting than the global average.
We computed the RCA index for each CPC/IPC category at a 1- and 3-digit level. Figure 3 shows the density distribution of the RCA index values by technological domains. While the distribution is centred below 1, indicating that green patents are on average more present in many non-specialised categories, where the left tail is slightly left-skewed. In other words, in general, there are green patents across many technological domains, whose main scope, however, is not to produce environmental technologies. When shifting to a more granular level of classification (please refer to Appendix B.2. Classification results, Tables B9, B10 and B11 for complete numbers of the RCA index at 3-digit of the CPC code), we observe how, specifically, top 3 positions are occupied by CPC Treatment of water, waste water, sewage, or sludge (C02), Technologies or applications for mitigation or adaptation against climate change (Y02), and Disposal of solid waste; reclamation of contaminated soil (B09).
Figure 3: Density plot of the RCA index by CPC category.
  • Note: The figure represents the density of the RCA index by CPC classification categories after applying our text-based analysis.
4.1.3 The dynamics of green patents
When we look at the dynamics of green patents according to previous literature in the period from 2010 to 2022, we observe an almost constant share of green over all patents in Figure 4, up to a point. The share is about 12% each year until the year 2022, when it peaks at $\simeq$ 21%. In contrast, when we leverage our custom definition of true green patents, the share remains almost unchanged at around 2.5% for the whole period considered, with the exception of a slight increase to 3.5% in 2022.
Figure 4: Granted green patents (grey) and true green patents (black) over total by publication year.
  • Note: The figure reports percentages of granted green patents (grey) and true green patents (black) over total granted applications by publication year (2010-2022).
Briefly, according to our evidence, there is no reason to observe a peak in green patents in 2021 and 2022. The increase is only in the technological classes that host environmentally relevant inventions. Once we look at the text of the supporting documents, the share of patents actually describing environmental technologies, i.e., what we call true greens, is not systematically changing over time.
4.2 Green citations
A previous exercise by (Shi et al., 2023) shows that a systematic portion of patents that are classified as green eventually do not receive many citations. In this section, we want to check whether this is the case for true green patent grants. We test the following basic equation on the subset of patents published in the technological classes that are classified as green by previous literature:
$\log(Y_{i(j)kt})=\beta_{0}+\beta_{1}True\_greens_{i(j)}+\beta_{2}X_{i(j)}+\delta_{ct}+\upsilon_{i(j)kt}$ (6)
where the dependent variable is the logarithm of the patent citations for each patent $i$ that belongs to a family $j$ in a technological class $k$ published in the priority year $t$. Please note that we transform the dependent variable by adding one before logs to avoid omitting the cases of citations equal to zero. The main coefficient of interest is the one on $True\_greens_{i(j)}$, which is a characteristic that is verified for each patent $i$, regardless of its technological class. In our baseline exercise, we add patent age and the number of patent family members in the matrix of controls $X_{i(j)}$. Finally, we add a battery of fixed effects to catch idiosyncrasies for expert evaluations in a technological class $k$ at priority year $t$.
Our baseline exercise is reported in column (4) of Table 4. We find that true greens are cited about $1\%$ less than the rest of the green patents, as from classifications in previous economic literature. Please note how the results change when we switch from column (2) to column (3) after we include fixed effects for technological classes and priority years. One possible explanation for the negative premium on citations by true green patents is that environmental technologies are relatively newer than brown technologies; therefore, there is less scope for citations by following novel inventions, which introduce other environmental technologies by relying on scarcer knowledge about environmental protection.
Table 4: True green citations premia
(1) (2) (3) (4)
True green -$0**$054*** -0.012*** -0.011***
(0.001) (0.001) (0.001) (0.001)
Controls No Yes Yes Yes
Class fe No No Yes No
Priority year fe No No Yes No
Class*priority fe No No No Yes
Family clusters Yes Yes Yes Yes
Adj. R squared 0.001 0.160 0.378 0.384
N. obs. 4,049,540 4,049,540 4,049,540 4,049,540
  • Note: The table reports results after dummy regressions where the dependent variable is the logarithm of the patent citations plus one. We consider the entire sample of green patents as considered by previous literature, and our coefficient of interest is the one on True green patents, after following our NLP approach. Patent controls include age and the number of patent family members. Errors are clustered at the patent family level. *** p$<$0.01, ** p$<$0.05, * p$<$0.1.
5 Green patenting and firms’ outcomes
In this second part of the paper, we focus on the relationship between green patenting and firms’ outcomes. We organise the work in the following way. First, we estimate statistical associations between patenting activity and a dashboard of financial accounts. Basic dummy regressions will show how much green and non-green patenting activity is associated with a battery of firms’ outcomes. Such a piece of preliminary empirical evidence lets us start forming an idea about the relationship between patenting and firm performance. Of course, positive associations do not imply any cause or effect because there is a complicated relationship between the innovation ability of a company, the protection of intellectual property rights it obtains, and the performance it records. Therefore, we discuss the endogeneity issues at the heart of the positive associations between patenting and firm performance. However, it is beyond the scope of this paper to fully unravel such endogeneity issues. Instead, we propose a simple identification strategy that solves at least the reverse causality problem between patenting and firms’ outcomes. Briefly, we know that larger, more productive, more profitable, and older firms are more likely to register a patent, and the positive associations cannot be informative of the gains from patenting; therefore, we want to purge this endogenous selection effect and obtain estimates of what a firm gains from patenting. Limitations are discussed at the end of this section. However, we believe that our estimates are valid to compare what happens when a company holds a true green patent, a non-true green patent, a non-green patent, and when it does not invest in patenting.
5.1 Firm-level premia
We perform basic dummy regressions on the dashboard of firms’ outcomes that we described in Section 3.2 after following the following basic equation:
$\ Y_{fsct}=\beta_{0}+\beta_{1}True\_greens_{ft}+\beta_{2}Patenting_{ft}+\beta_{2}size_{ft}+\gamma_{s}+\theta_{c}+\lambda_{t}+\upsilon_{fsct}$ (7)
where on the left-hand side we separately test sales, market share, labor productivity, capital intensity, ROCE, EBIT, and TFP. We perform basic least-squares with firm-level clustered errors when the dependent variable is transformed into the logarithm of firm-level outcomes. The only exception is the ROCE, which can have negative values, and it cannot be transformed into a logarithm. For the latter, we adopt the inverse hyperbolic sine transformation proposed by Bellemare and Wichman (2020). The two main coefficients of interest are True greens and Patenting. True greens is equal to one when the firm has been granted at least one patent that has been acknowledged by our algorithm to be green in that year, and zero otherwise. Patenting is equal to one if the firm has been granted at least one patent regardless of the technological classes, and zero otherwise. Fixed effects are included by sector, country, and time.
Table 5: True green patent grants and firms’ outcomes - correlations
VARIABLES Sales Mkt share Lab. prod. K-intensity ROCE EBIT TFP
True greens 0.253 0.036*** 0.072** 0.172*** -0.213*** -1.149*** -0.101
(0.192) (0.008) (0.034) (0.028) (0.70) (0.420) (0.124)
Patenting 1.957*** 0.013*** 0.345*** 0.204*** -0.063*** 0.608*** 0.085***
(0.027) (0.001) (0.009) (0.007) (0.018) (0.112) (0.030)
Controls:
Firm size
2-digit industry fe
Country fe
Year fe
N. firms with green grants 3,137 3,137 2,407 3,111 3,137 2,576 1,133
N.obs. 11,320,979 11,320,979 9,337,060 11,253,885 11,320,984 8,997,862 1,784,972
Adj. R-squared 0.938 0.387 0.951 0.222 0.420 0.187 0.870
  • Note: The table reports results after dummy regressions where dependent variables are a chosen dashboard of firms’ outcomes, and the coefficients of interest are True green and Patenting. Firm size is measured by the number of employees, and industries are 2-digit NACE rev. 2. Sales, labor productivity, capital intensity, and TFP are expressed in logarithms. ROCE can have zero values and is transformed with a hyperbolic sine as proposed by Bellemare and Wichman (2020). Market share and EBIT are expressed in levels. Firm-level clustered standard errors. *** p$<$0.01, ** p$<$0.05, * p$<$0.1.
In particular, we observe that, after controlling for our NLP approach, firms that obtain true green show positive premia in market shares, labor productivity, and capital intensity. Interestingly, we record negative premia on both measures of profitability, the ROCE and the EBIT, while the coefficient of TFP for manufacturing firms is not statistically significant. Please note how important it is to control for Patenting. Regardless of the technology, firms that obtain a patent grant show almost always higher outcomes, with the notable exception of profitability measures.
At this stage, we cannot interpret the coefficients in Table 5 as the effect of patenting activity on firms’ competitiveness. We can only conclude that, apparently, firms with true green grants are statistically different from the ones with grants in other technological classes. In the next Section, we will provide a few additional elements to comment on thier statistical differences.
5.2 Green patenting, firms’ outcomes, and reverse causality
In this Section, we want to specifically address the endogenous selection of firms into true green patenting. Given previous positive associations between true green grants and a few firms’ outcomes, it is quite possible that a large part of that statistical association is due to the ability of bigger and more productive firms to invest in green technologies and obtain a patent grant. It is a textbook case of reverse causality, as the direction of cause and effect is unclear and potentially inverted.
To challenge reverse causality, we design a basic identification strategy with a propensity score matching (PSM). Our aim is to simulate a randomised experimental setting by pairing firms with similar observable characteristics (like industry, country, firm size, past performance) where one receives the treatment, in our case a green grant, and the other does not. Therefore, by comparing the firms that are matched, we isolate the impact of the green grants on firms’ performance, possibly eliminating the bias from reverse causality. Please note that we believe that our albeit simple PSM strategy is powerful in our case, because it allows us to control for observable characteristics that make firms more likely to get a patent grant.
Table 6 shows the distribution of treated firms in our panel. The number of newly treated individuals ranges from 372 to 241 per year. Please note that the group of firms that have never obtained a true green grant is huge, and that is the set from which we need to obtain a counterfactual after our PSM.
Table 6: Distribution of treated individuals in our panel over time.
Year True green
2010 329
2011 372
2012 359
2013 347
2014 352
2015 327
2016 335
2017 369
2018 348
2019 373
2020 353
2021 318
2022 241
  • Note: The table reports the panel distribution of companies in our sample that had at least one true green grant, i.e., following our NLP classification, in our period of analysis. Please note that repetitions may occur if the same firm had a green patent granted in different years, this definition of treatment is in line with our cross-section application of the PSM model.
Propensity scores are obtained after a logistic regression where we introduce the following covariates: i) the first lag of capital intensity in logs, ii) year-over-year variation of capital intensity in logs, iii) the first lag of the ROCE, iv) year-over-year variation of the ROCE, v) the firm’s age in logs, and, finally, vi) a battery of country, NACE 2-digit, and year fixed effects. After obtaining the scores, we pick the nearest neighbour without replacement and, to improve the quality of the matches, we impose common support on the distribution of covariates both in the treated and the control group.
After the match of each firm with at least one true green grant with its nearest-neighbour that did not, we can estimate the Average Treatment Effect on the Treated (ATET) by comparing the means of the two distributions. In Table 7, we find that statistically significant ATETs emerge primarily for sales, market shares, and labor productivity. If a company obtains a true green grant, it can rely on a remarkable increase in sales ($e^{1.357-1}\approx$ 143%), market share ($e^{0.033-1}\approx$ 38%), and labor productivity ($e^{0.085-1}\approx$ 40%). These effects align with the correlations observed in previous paragraphs. The main departure lies in the coefficients of capital intensity, ROCE, EBIT, and TFP. Profitability measures are not statistically significant after we introduce a matching strategy, while they correlate negatively with true green grants. The most likely explanation is that, after investing in green technologies, firms are, on average, at a loss. They still need to break even after the sunk cost of the R&D needed to obtain those grants. Yet, after we construct our counterfactual, there is no systematic difference in profitability between a company that invests in green technologies and a similar one that invests in any other technology.
In summary our NLP approach to identifying green patents appears to better capture the impact of environmental technologies on top of the gains that companies would have in any case from patenting.
Table 7: Average Treatment Effects on the Treated after a propensity score matching - True green patents and firms’ outcomes
Sales Mkt share Lab. prod. K-intensity ROCE EBIT TFP
ATET 1.357*** 0.033*** 0.085*** -0.01 0.048 0.11 0.002
std. error 0.089 0.007 0.028 0.027 0.073 0.418 0.156
t-stat. 15.232 4.782 3.053 -0.372 0.655 0.263 0.014
Treated obs. 2,615 2,276 2,034 2,276 2,276 2,076 968
Untreated obs. 46,345 41,276 36,378 41,276 41,276 37,432 24,992
  • Note: The table reports ATTs computed via PSM. The treated companies have at least one true-green patent following our NLP approach in a single year, and it is matched with companies that have been granted at least one patent in 2010-2022, regardless of the technology nature. Propensity score matching is performed by nearest-neighbour without replacement, assuming a common support on the covariates. *** p$<$0.01, ** p$<$0.05, * p$<$0.1.
6 Robustness and sensitivity checks
In this section we present robustness and sensitivity controls on previous results using our measures. The inclusion of the measure following Arts et al. (2021) reveals that the positive effects of green patenting are highly contingent on the novelty they introduce in their technology classes. Our NLP approach, used together with the NLP approach for novelty (8), suggests that green patenting delivers stronger economic returns when it is genuinely transformative.
Table 8: Average Treated Effects on the Treated after a propensity score matching - high-novelty true-green patent grants vs high-novelty patents.
Sales Mkt share Lab. prod. K-intensity ROCE EBIT TFP
ATET 0.885*** 0.026 -0.014 -0.086 -0.047 0.409 0.403
std. error 0.243 0.018 0.067 0.058 0.142 0.875 0.464
t-stat. 3.643 1.433 -0.217 -1.481 -0.33 0.468 0.87
Treated obs. 530 499 440 499 499 467 157
Untreated obs. 11,135 10,421 9,331 10,421 10,421 9,841 5,724
  • Note: The table reports ATETs computed after a propensity score matching. The treated unit is a company that has been granted at least one high-novelty true-green patent for a single year, and it is matched with companies that have at least one high-novelty grant, regardless of the technology nature, in the period 2010-2022. High novelty is defined following the NLP approach by Arts et al. (2021), while true greens are defined following our NLP approach. Propensity score matching is performed by nearest-neighbour without replacement, assuming a common support on the covariates. *** p$<$0.01, ** p$<$0.05, * p$<$0.1
As a first robustness check, we estimate the ATETs separately for small and large firms (Tables 10 and 9). Firms are classified based on the median number of employees [^9] Measured as the average number of employees over the 2010–2022 sample period in our sample, which is 3.6. For larger firms, the results are largely consistent with the baseline estimates. In contrast, for smaller firms, no statistically significant effects are observed across most outcomes, with the exception of a 14% positive effect on EBIT. This result may suggest that while smaller firms do not immediately experience growth in scale (e.g., sales or market share), they may achieve gains in operating efficiency or cost reductions from engaging in green patenting activities. In contrast, larger firms may already benefit from economies of scale and more established innovation pipelines, making the marginal benefit of green patenting less pronounced in profitability terms.
Table 9: Average Treated Effects on the Treated after a propensity score matching - true-green patent and firms outcomes for large firms.
Sales Mkt share Lab. prod. K-intensity ROCE EBIT TFP
ATET 1.382*** 0.032*** 0.093*** -0.046* 0.07 0.058 0.006
std. error 0.085 0.007 0.027 0.027 0.072 0.394 0.158
t-stat. 16.198 4.634 3.465 -1.727 0.967 0.148 0.036
p-value 0.0 0.0 0.001 0.084 0.333 0.883 0.971
Treated obs. 2,573 2,247 2,009 2,247 2,247 2,050 962
Untreated obs. 45,124 40,313 35,729 40,313 40,313 36,622 24,784
  • Note: The table reports ATETs computed after a propensity score matching only for firms with an average size (determined by number of employees) during the sample period, above the median value. The treated unit is a company that one true-green patent following our NLP approach in a single year, and it is matched with companies that have been granted at least one patent in 2010-2022, regardless of the technology nature. Propensity score matching is performed by nearest-neighbour without replacement, assuming a common support on the covariates. *** p$<$0.01, ** p$<$0.05, * p$<$0.1
Table 10: Average Treated Effects on the Treated after a propensity score matching - true-green patent and firms outcomes for small firms.
Sales Mkt share Lab. prod. K-intensity ROCE EBIT TFP
ATET -0.353 0.0 -0.078 -0.235 0.313 14.714** 0.187
std. error 0.72 0.003 0.509 0.452 0.881 7.349 0.217
t-stat. -0.49 0.057 -0.154 -0.519 0.355 2.002 0.862
p-value 0.624 0.954 0.878 0.604 0.723 0.046 0.395
Treated obs. 42 28 25 28 28 25 6
Untreated obs. 727 508 366 508 508 440 29
  • Note: The table reports ATETs computed after a propensity score matching only for firms with an average size (determined by number of employees) during the sample period, below or equal the median value. The treated unit is a company that one true-green patent following our NLP approach in a single year, and it is matched with companies that have been granted at least one patent in 2010-2022, regardless of the technology nature. Propensity score matching is performed by nearest-neighbour without replacement, assuming a common support on the covariates. *** p$<$0.01, ** p$<$0.05, * p$<$0.1
A second robustness check involves distinguishing between firms located in “old” and “new” EU member states (Tables 11 and 12). Specifically, we define old members as countries that joined the European Union before 2000, and new members as those that joined thereafter. No significant ATETs are found for firms in the new member states. However, for firms in the old member states, we observe a positive and statistically significant effect on both sales ($e^{0.803-1}\approx$ 82%) and market share ($e^{0.05-1}\approx$ 39%). One possible interpretation is that firms in old EU member states are better positioned to capitalize on green patenting activities due to more mature institutional environments, stronger regulatory frameworks, and better access to green innovation subsidies or infrastructure. In contrast, firms in new member states may still be in the process of integrating environmental standards and innovation systems, limiting their ability to convert green patenting into immediate competitive or market advantages. Alternatively, differences in domestic demand for green products and varying levels of international exposure could also contribute to the observed heterogeneity in outcomes.
Table 11: Average Treated Effects on the Treated after a propensity score matching - true-green patent and firms outcomes for first countries joining EU.
Sales Mkt share Lab. prod. K-intensity ROCE EBIT TFP
ATET 0.803*** 0.05** 0.035 -0.02 0.027 0.039 -0.279
std. error 0.261 0.02 0.054 0.056 0.165 0.88 0.39
t-stat. 3.078 2.544 0.652 -0.367 0.163 0.044 -0.715
Treated obs. 457 432 378 432 432 403 137
Untreated obs. 1,724 1,579 1,421 1,579 1,579 1,424 720
  • Note: The table reports ATETs computed after a propensity score matching only for firms of countries joining the EU before the year 2000, i.e. Austria, Belgium, Denmark, Finland, France, Germany, Greece, Ireland, Italy, Luxembourg, Netherlands, Portugal, Spain and Sweden. The treated unit is a company that one true-green patent following our NLP approach in a single year, and it is matched with companies that have been granted at least one patent in 2010-2022, regardless of the technology nature. Propensity score matching is performed by nearest-neighbour without replacement, assuming a common support on the covariates. *** p$<$0.01, ** p$<$0.05, * p$<$0.1
Table 12: Average Treated Effects on the Treated after a propensity score matching - true-green patent and firms outcomes for last countries joining EU.
Sales Mkt share Lab. prod. K-intensity ROCE EBIT TFP
ATET -0.141 0.043 0.119 -0.071 -0.177 3.31 -2.377
std. error 0.454 0.044 0.132 0.157 0.366 2.889 3.497
t-stat. -0.31 0.965 0.908 -0.452 -0.483 1.146 -0.68
Treated obs. 61 54 51 54 54 51 14
Untreated obs. 247 115 101 115 115 110 37
  • Note: The table reports ATETs computed after a propensity score matching only for firms of countries joining the EU after the year 2000, i.e. Bulgaria, Croatia, Cyprus, Czechia, Estonia, Hungary, Latvia, Lithuania, Malta, Poland, Romania, Slovakia and Slovenia. The treated unit is a company that one true-green patent following our NLP approach in a single year, and it is matched with companies that have been granted at least one patent in 2010-2022, regardless of the technology nature. Propensity score matching is performed by nearest-neighbour without replacement, assuming a common support on the covariates. *** p$<$0.01, ** p$<$0.05, * p$<$0.1
7 Conclusions and final remarks
This study’s primary contribution is the development and application of an NLP methodology for classifying “true” green patents. Our research suggests that this custom NLP approach can refine the identification of environmentally relevant patents compared to traditional, broader classification systems. Standard methods may risk conflating genuinely environmental technologies with conventional technologies. Our validation procedure, comparing the NLP approach against more conventional classifications, indicates that standard methods broadly overestimate green patents. We find that about 20% of those otherwise flagged as green according to standards in literature are actually true-green patents when we apply our text-analysis methodology. We show, however, an important degree of heterogeneity once we scroll down the different technological classes that are touched upon by true environmental technologies.
Notably, in the second part of the paper, we show that companies that have been granted at least one true green patent show remarkable increases in sales, market shares, and labor productivity. Although we are not able to separate the impact of IPR protection from the implicit innovative ability of the patenting firm, we believe that our findings are useful to discuss the premia that companies obtain.
Importantly, in this contribution, we show that it is possible to look into supporting documents and systematically classify patents by means of text analysis, going beyond standard technological classes. Our custom NLP classification framework offers a methodological approach that could be adapted to other domains. There are many contexts in which a custom classification may assist policymakers and researchers in gauging the impact of patenting activity.
Appendix A: Natural language processing
A.1. Green-patents vocabulary
This section contains reference to our custom dictionary constructed for true green patents classification. Table A1 contains baseline expressions related to green patents and technologies produced by Chat GPT and Gemini and used to feed the Word2Vec model for the production of an enlarged dictionary. Table A2, A3, and A4 represent our final enlarged vocabulary for regex match into patents descriptive texts. It is made of two kinds of expressions:
  • Single expressions like energy conservation that, if matched into a text, would label it as green related.
  • Co-occurrences like carbon and capture that would require the presence of both at a minimum distance of 20 words to classify a patent text as green.
Table A1: Baseline key-expresions.
Baseline keywords
alternative energy waste management energy recovery ozone generators
anaerobic digestion waste reduction energy savings photocatalysis
carbon capture water conservation engine optimization plant based
carbon footprint aerodynamic design environmental certification pollution control
carbon sequestration air purification environmental compliance precision agriculture
circular economy bio-based materials environmental conservation product lifespan
clean energy biodegradable polymers environmental labeling public transportation
clean production biodegradable products environmental management rainwater harvesting
composting biodiversity environmental policy recyclable materials
conservation biofuel environmental stewardship recycled materials
dioxide capture biomass fuel cells recycled packaging
energy conservation bioplastics geothermal recycled products
energy efficient biotechnology green buildings resource efficient
energy optimization carbon filters green design smart grids
environmental impact carbon neutral green innovation solar geoengineering
environmental risk carbon offsetting green manufacturing solar panels
geothermal power cellulosic plastics green technology sustainability metrics
global warming clean technology green transportation sustainability practices
green chemistry climate adaptation greenhouse gas sustainable agriculture
hydropower climate friendly greenhouse gases sustainable farming
low carbon compostable materials hybrid vehicles sustainable forestry
low emission eco conscious hydroelectric dams sustainable manufacturing
membrane filtration eco efficient hydroponics sustainable materials
membrane separation eco friendly insulation sustainable practices
recycling eco innovation leed certified vertical farming
regenerative braking eco solutions life cycle waste minimization
renewable energy ecodesign lightweight materials waste recycling
resource recovery ecological design low impact water efficient
reverse osmosis ecological footprint material compatibility water treatment
sustainable electric motors natural ventilation wind turbines
sustainable development electric vehicles organic farming zero emissions
urban planning energy efficiency ozone depletion zero waste
  • Note: The table reports the list of baseline key-expresions related to green technologies generated by ChatGPT and Gemini. From this baseline we performed vocabulary expansion with the use of Word2Vec.
Table A2: Keywords for single REGEX match
Expressions for single match
aerogenerator carbon filters ecological design geothermal photocatalysis soilless
aerodynamic design carbon neutral ecological footprint geothermal power photocatalyst soilless cultivation
aeroponic carbon offset ecological restoration geothermy photocatalytic soilless culture
afforestation carbon offsetting ecologically global warming photodegradable solar geoengineering
air purification carbon sequestration ecology green buildings photolysis solar panels
alternative energy carbonfilter ecosystem green chemistry photolytic solarenergy
anaerobic catalyzing electric motors green design photooxidation solarpanel
anaerobic digestion cellulosic plastics electric powertrain green innovation photovoltaic solarpower
bioactivity circular economy electric vehicle green manufacturing plant based sustainability
bio-based materials clean energy electric vehicles green technology pollution control sustainability metrics
biocarbon clean production electricvehicle green transportation polylactide sustainability practices
biochemistry clean technology emission reduction greenhouse gas precision agriculture sustainable
biocomposite climate adaptation energy conservation greenhouse gases product lifespan sustainable agriculture
biodegradable climate friendly energy efficiency greening public transportation sustainable development
biodegradable compostable compostable energy efficient hybrid powertrain rainwater harvesting sustainable farming
biodegradable polymer compostable material energy optimization hybrid vehicle recyclable sustainable forestry
biodegradable polymers compostable materials energy recovery hybrid vehicles recyclable material sustainable manufacturing
biodegradable product composted energy save hydroelectric recyclable materials sustainable materials
biodegradable products composting energy saving hydroelectric dam recyclate sustainable practices
biodegrade conservancy hydropower energy savings hydroelectric dams recycle material vertical farming
biodiesel conservation energysaving hydroponics recycle practice thrift waste minimization
biodiversity decarbonation engine optimization hydropower recycle product waste recycle
bioenergy decarbonization environmental certification insulation recycleable waste recycling
bioethanol dechlorinate environmental compliance leed certified recycled waste reduction
biofertilizer decyanation environmental conservation life cycle recycled materials water conservation
biofuel demercuration environmental labeling lightweight materials recycled packaging water efficient
bioleache depollute environmental management lignocellulose recycled products water purification
biologically degradable depollution environmental policy low emission recycling water treatment
biomass development environmental stewardship low impact regenerative braking waterpower
bioplastic dioxide capture environmentally friendly mangrove forest renewable energy wind turbines
bioplastics eco conscious environmentally sustainable material compatibility resource efficient windmill
biopolymer eco efficient fertigation natural ventilation resource recovery windpower
bioproduct eco friendly footprint accounting organic farming rooftop greening windturbine
bioremediation eco innovation forward osmosis oxyfuel seawater desalination windwheel
biotechnology eco solutions fuel cells ozone depletion seawater desalting zero emissions
carbon capture ecodesign fuelcell ozone generators smart grids zero waste
  • Note: The table reports the list of keywords for single REGEX match into patents texts, i.e. a text containing at least one of the expression, if classified as “green” by code-based algorithms, were classified as “true green”.
Table A3: Keywords combinations for co-occurrence REGEX match (part I).
Keyword Co-occurrence
activatedcarbon Filtration OR purification OR pollutants
aeroponic cultivation Hydroponics OR vertical farming
aeroponic culture Hydroponics OR vertical farming
bamboo charcoal Sustainable forestry
biochemical Biofuels OR bioplastics
bioconversion Biomass OR renewable energy
butanol Biofuels
carbon Capture
carwash Water-efficient
claus Sulfur recovery OR tail gas treatment
daylighting energy saving
deduste Air pollution
denitride Air pollution
denitrifying Water pollution
denitrogenation Air pollution
denox Air pollution
desalinate Energy-efficient
desalination Energy-efficient
desalinization Energy-efficient
desalt Energy-efficient
desulfurate Air pollution
disinfection Non-chlorine OR ultraviolet
disinfector Non-chlorine OR ultraviolet
dmfc Direct methanol fuel cell
electrochemical Fuel cell
electrodialysis Desalination OR wastewater treatment
electrolyzer Hydrogen production OR water electrolysis
environmental Sustainability OR pollution reduction
enzymolysis Biodegradable materials OR biomass conversion
ethanol Biofuel OR renewable energy
evaporative crystallization Water treatment OR desalination
fenton Wastewater treatment OR pollutant degradation
ferment Biofuels OR biogas production
fermentable sugar Biofuels OR biogas production
filter Air pollution control OR water treatment
flywheel Energy storage OR regenerative braking
greywater Water reuse OR wastewater treatment
humification Composting ganic matter
hydrolysed Biomass conversion OR biofuel production
hydrolytic acidification Biomass conversion OR biofuel production
hyperfiltration Water treatment OR desalination
insulate Energy efficiency OR building insulation
insulated Energy efficiency OR building insulation
insulating Energy efficiency OR building insulation
insulative Energy efficiency OR building insulation
insulator Energy efficiency OR building insulation
lignocellulosic Biomass conversion OR biofuel production
livestock excreta Biogas production OR waste management
livestock manure Biogas production OR waste management
membrane Filtration OR water treatment
methanation Biogas production OR renewable energy
methanogenesis Biogas production OR renewable energy
microfiltration Water treatment OR wastewater treatment
microgrid Renewable energy OR distributed generation
nanobubble Water treatment OR air purification
nanofilter Water treatment OR wastewater treatment
nanofiltration Water treatment OR wastewater treatment
nitrifying denitrify Wastewater treatment OR nitrogen removal
oilfield reinjection Enhanced oil recovery OR carbon capture
orc Organic Rankine Cycle OR waste heat recovery
cropping practice Sustainable OR organic
energysource Clean OR sustainable
foodwaste Reduction OR avoid
renewable Energy OR resource
  • Note: The table reports the list of keywords co-occurrences for REGEX match into patents texts, i.e. a text containing at least one of the expression combinations, if classified as “green” by code-based algorithms, were classified as “true green”.
Table A4: Keywords combinations for co-occurrence REGEX match (part II)
Keyword Co-occurrence
renewably Energy OR resource
reusable Material OR product OR resource
reuse Material OR product OR resource
ordered mesoporous catalysis
oxygenator Aquaculture OR wastewater treatment
oxygenerator Aquaculture OR wastewater treatment
ozonator Water treatment OR air purification
ozone Water treatment OR air purification
ozonize Water treatment OR air purification
ozonizer Water treatment OR air purification
ozonolysis Water treatment OR air purification
peak shaving energy management OR energy storage
pemfc Proton exchange membrane fuel cell OR hydrogen fuel cell
pervaporation Membrane separation OR desalination
photochemistry Solar energy conversion
photoreaction Solar energy conversion
photosynthetic Biofuels
poultry farming Sustainable agriculture OR waste management
poultry manure Sustainable agriculture OR waste management
propeller Electric OR hybrid OR low-emission
propulsion Electric OR hybrid OR low-emission
propulsor Electric OR hybrid OR low-emission
pyrohydrolysis Waste conversion OR biomass conversion
recuperation energy OR waste OR material
regasification Liquefied natural gas OR renewable energy sources
regenerative Braking OR energy recovery
remediation Environmental cleanup OR contaminated soil
reprocess Waste management OR resource recovery
reprocessed Waste management OR resource recovery
reprocessing Waste management OR resource recovery
resource Efficiency OR conservation
restoration Ecological restoration OR land reclamation
reusing Waste OR Material
reutilization Waste OR Material
reutilize Waste OR Material
reutilizing Waste OR Material
reverseosmosis Water desalination OR water treatment
rice husk Biomass conversion OR renewable energy source
saccharify enzyme Biofuel production OR biomass conversion
seedling raising Sustainable agriculture OR forestry
semicoke Biomass conversion OR coal alternative
simultaneously practice thrift Resource efficiency OR waste reduction
smart grid Renewable energy integration OR energy efficiency
sncr denitration Flue gas treatment OR air pollution control
sofc Solid oxide fuel cell OR clean energy
solar Energy OR cell OR panel
straw Biomass resource OR renewable energy
sugar beet Biofuel feedstock OR renewable energy
sun shading Sustainable design OR building efficiency
sunshade Sustainable design OR building efficiency
supercapacitor Energy storage OR renewable energy integration
supercharge Electric vehicle OR battery
supercharging Electric vehicle OR battery
thermocatalytic Biomass conversion OR waste valorization
thermochemical Biomass conversion OR renewable energy
ultrafilter Water treatment OR purification
ultrafiltrate Water treatment OR purification
ultrafiltration Water treatment OR purification
ultralight Material OR construction
ultralow Emission OR power consumption
uv photolysis Water treatment OR pollutant degradation
wasted Recycling OR resource recovery
wastepaper Recycling OR paper recycling
  • Note: The table reports the list of keywords co-occurrences for REGEX match into patents texts, i.e. a text containing at least one of the expression combinations, if classified as “green” by code-based algorithms, were classified as “true green”.
A.2. Assessing patent novelty
In this section, we describe the methodologies employed to evaluate the novelty of patents to better understand their potential economic impact. Specifically, we followed the methodology of Arts et al. (2021), and implemented a series of steps to quantify the novelty and technological significance of each patent, with the production of a number of metrics:
  • Base dictionary construction: We constructed a base dictionary using the text of all patent abstracts filed before 1980 (the whole PATSTAT Global database). This historical corpus provided a benchmark against which we could measure the novelty of more recent patents.
  • Unique word analysis: For each patent in our dataset, we computed the number of unique words, as well as pairs and triplets of consecutive words, that appeared for the first time in the patent literature. This measure helped us gauge the linguistic and conceptual innovation embodied in each patent.
  • Pairwise keyword combinations: Finally, we computed the number of unique combinations of keywords that had not previously appeared together in earlier patents, further capturing the innovative potential of the patent.
In their paper, the authors did not identify a strategy for the simultaneous use of these metrics but compared the results between them (and between classical metrics of innovation content of patents). Since the new pairwise keyword combinations metric was the one better performing in identifying highly innovative patents according to the authors, we used such index as a proxy of the novelty of an application for our analysis.
A.3. Word2Vec specification
In this section, we provide details about the selection of the Word2Vec model’s hyperparameters. We need to find the context window size ($C$), which defines the context length for training the neural network, and the minimum count ($MC$), which sets the frequency threshold for a word to be included in the training corpus. We employ a data-driven tuning process. We use WordSim353, which is a test collection for measuring word similarity or relatedness (Agirre et al., 2009). We then train multiple Word2Vec models using a grid of different $W$ and $MC$ parameter combinations. For each trained model, we calculated the cosine similarity for word pairs in the WordSim353 list and measured the Pearson correlation between our results and the gold standard scores. As shown in Figure A1, all tested parameter pairs yielded a positive correlation above 50%. Generally, models with a smaller $C$ performed better when the minimum count was held constant. This empirical process pointed towards optimal values of approximately $MC$=40 and $C$=2, which we adopt in our analysis in the manuscript.
Figure A1: Grid search for the selection of context window and minimum count hyperparameters.
  • Note: The figure represents the hyperparameters’ selection for the Word2Vec model. Each point corresponds to a combination of parameters, and the related linear correlation value with our benchmark similarity standard (the higher the better).
We use the same procedure for fine tuning the embedding vector size $d$ (first layer of our shallow neural network). Figure A2 show correlation with our benchmark for different sizes of embeddings. Given the convergence around 0.61 for embedding with size greater than 450, we choose such value in order to get the best empirical performances while minimizing the model complexity and overcoming limitations on computational power in larger settings.
Figure A2: Grid search for the selection of the embedding vector size.
  • Note: The figure represents the selection of the embedding dimension for the Word2Vec model. Each point corresponds to a value of vector size, and the related linear correlation value with our benchmark similarity standard (the higher the better).
Appendix B: Tables and graphs
B.1. Firm-level dataset distribution
In this section, we describe the distribution of firm-level data used for our econometric application. Table B1 represent the distribution of firms over European countries, with the majority of firms located in Italy (20.93%), Spain (14.97%) and France (8.57%).
Table B1: Distribution of firms over countries.
Country N. Firms % of total
Austria 2,813 0.11
Belgium 21,436 0.80
Bulgaria 184,784 6.90
Croatia 58,882 2.20
Cyprus 284 0.01
Czechia 25,284 0.94
Denmark 92 0.01
Estonia 63,579 2.38
Finland 90,140 3.37
France 229,293 8.57
Germany 13,550 0.51
Greece 3,556 0.13
Hungary 193,670 7.24
Ireland 44 0.01
Italy 560,255 20.93
Latvia 35,433 1.32
Lithuania 4,782 0.18
Luxembourg 439 0.02
Malta 1,585 0.06
Netherlands 2,017 0.08
Poland 64,232 2.40
Portugal 178,350 6.66
Romania 234,546 8.76
Slovakia 35,988 1.34
Slovenia 68,132 2.55
Spain 400,753 14.97
Sweden 202,596 7.57
  • Note: The table reports the distribution of firms of our dataset over countries. Data are related to corporations in Europe with available financial data over the considered period.
Table B2 represent the distribution of firms by size proxied by number of employees. The majority of firms in our sample (62.76%) have less than 20 employees, while the second largest category is represented by firms with more than 1,000 employees, accounting for 26.47%. Table B3 represent the distribution of firms by 2-digits NACE. Sectors wholesale and retail trade (45-47), manufacture (10-32), and financial, insurance and real estate activities (64-68) together account for almost 50% of our sample, with the remaining firms distributed among other NACE accounting for less than 5% each.
Table B2: Distribution of firms by size.
Size N. Firms % of total
1-20 2,129,110 62.76
21-250 292,139 9.83
251-1000 26,714 0.94
>1000 542,105 26.47
  • Note: The table reports the distribution of firms of our dataset by size (i.e. n. of employees). Data are related to corporations in Europe with available financial data over the considered period - snapshot of year 2018.
Table B3: Distribution of firms by 2-digits NACE.
NACE Description N. Firms % of total
01-03 Crop, animal production, fishing, hunting, forestry and logging 79,028 3.16
05-09 Mining and extraction 5,914 0.24
10-32 Manifacture 313,035 12.5
33 Repair and installation of machinery and equipment 22,318 0.89
35 Electricity, gas, steam and air conditioning supply 24,245 0.97
36-39 Water, sewerage and waste management 15,266 0.61
41-43 Construction, civil engineering and special activities 125,066 4.99
45-47 Wholesale and retail trade 630,142 25.16
49-51 Transport 97,219 3.88
52 Warehousing and support activities for transportation 26,388 1.05
53 Postal and courier activities 2,514 0.10
55 Accommodation 40,433 1.61
56 Food and beverage service activities 92,873 3.71
58 Publishing activities 14,375 0.57
59-63 ICT, motion pictures, television and broadcasting 96,546 3.86
64-68 Financial, insurance and real estate activities 284,155 11.35
69 Legal and accounting activities 76,168 3.04
70 Activities of head offices; management and consultancy 81,255 3.24
71 Architectural and engineering activities 76,801 3.07
72 Scientific research and development 6,824 0.27
73 Advertising and market research 28,086 1.12
74 Other professional, scientific and technical activities 37,452 1.50
75 Veterinary activities 5,314 0.21
77 Rental and leasing activities 18,562 0.74
78 Employment activities 6,942 0.28
79 Travel agency and tour operator 14,402 0.58
80 Security and investigation activities 6,712 0.27
81 Services to buildings and landscape activities 27,566 1.10
82 Office administrative and office support 34,011 1.36
84 Public administration and defence 511 0.02
85 Education 32,992 1.32
86 Human health activities 74,445 2.97
87 Residential care activities 7,731 0.31
88 Social work activities without accommodation 7,793 0.31
90-91 Cultural and creative activities 13,935 0.56
92-93 Sports, amusement and gambling activities 25,394 1.01
94 Activities of membership organisations 1,340 0.05
95 Repair of computers and personal goods 10,743 0.43
96 Other personal service activities 39,633 1.58
97 Households as employers ofdomestic personnel 23 0.01
98 Undifferentiated goods andservices 7 0.01
99 Activities of extraterritorial organisations and bodies 51 0.01
  • Note: The table reports the distribution of firms of our dataset by 2-digits NACE. Data are related to corporations in Europe with available financial data over the considered period.
Table B4 represent the distribution of innovating firms in our sample, by country. Italy, France and Germany are the countries hosting the largest number of firms innovating regardless of the domain/definition. Table B5 shows the same distribution over 2-digits NACE. Also on this level the ranking remains the same regardless of the innovation domain or definition (i.e. green vs. true green), with sectors manufacture (10-32), wholesale and retail trade (36-39) and scientific research and development (72) occupying the top 3 positions.
Table B4: Distribution of innovating firms by country.
Country Patent Green True green
Austria 385 147 28
Belgium 529 179 44
Bulgaria 206 18 2
Cyprus 1 0 0
Czechia 626 121 26
Germany 2,104 743 168
Denmark 31 16 7
Estonia 44 9 3
Spain 1,200 232 41
Finland 813 213 44
France 2,681 668 170
Greece 16 3 1
Croatia 20 3 1
Hungary 108 25 4
Ireland 6 4 1
Italy 3,852 626 111
Lithuania 19 5 1
Luxembourg 22 9 2
Latvia 7 1 0
Malta 7 2 0
Netherlands 187 63 17
Poland 984 228 66
Portugal 125 30 4
Romania 146 44 20
Sweden 1,156 271 33
Slovenia 86 19 1
Slovakia 101 21 1
  • Note: The table reports the distribution of firms of our dataset with at least one patent, green patent or true green patent, by country. Data are related to corporations in Europe with available financial data over the considered period.
Table B5: Distribution of innovating firms by 2-digits NACE.
NACE Description Patent Green True green
01-03 Crop, animal production, fishing, hunting, forestry and logging 56 12 3
05-09 Mining and extraction 59 31 9
10-32 Manifacture 9,602 2,110 426
33 Repair and installation of machinery and equipment 169 39 4
35 Electricity, gas, steam and air conditioning supply 96 70 19
36-39 Water, sewerage and waste management 100 69 19
41-43 Construction, civil engineering and special activities 388 118 17
45-47 Wholesale and retail trade 1,520 275 50
49-51 Transport 35 10 1
52 Warehousing and support activities for transportation 40 9 0
53 Postal and courier activities 6 2 1
55 Accommodation 3 0 0
56 Food and beverage service activities 12 2 0
58 Publishing activities 67 3 1
59-63 ICT, motion pictures, television and broadcasting 383 51 7
64-68 Financial, insurance and real estate activities 391 87 17
69 Legal and accounting activities 21 8 3
70 Activities of head offices; management and consultancy 434 141 33
71 Architectural and engineering activities 636 207 54
72 Scientific research and development 874 300 84
73 Advertising and market research 24 1 0
74 Other professional, scientific and technical activities 115 34 8
75 Veterinary activities 2 0 0
77 Rental and leasing activities 78 12 6
78 Employment activities 7 2 2
79 Travel agency and tour operator 3 1 1
80 Security and investigation activities 17 3 0
81 Services to buildings and landscape activities 17 4 0
82 Office administrative and office support 76 22 3
84 Public administration and defence 4 1 0
85 Education 92 56 26
86 Human health activities 82 12 0
87 Residential care activities 4 0 0
88 Social work activities without accommodation 2 1 0
90-91 Cultural and creative activities 3 0 0
92-93 Sports, amusement and gambling activities 12 1 1
94 Activities of membership organisations 7 4 0
95 Repair of computers and personal goods 8 0 0
96 Other personal service activities 17 2 1
  • Note: The table reports the distribution of firms of our dataset with at least one patent, green patent or true green patent, by 2-digits NACE. Data are related to corporations in Europe with available financial data over the considered period.
B.2. Classification results
In this section, we provide descriptive statistics about both kind of green patents classification implemented in this analysis, i.e. green and true green. Specifically, Tables B6, B7 and B8 contain number of patents, green patents and true green patents granted in every CPC/IPC technological domain (at a 3-digits level). Tables B9, B10 and B11 represent our RCA index computed for each CPC classes again at a 3-digit level definition.
Table B6: Patents distribution by CPC (part I).
Classification code Description Patents Green patents True green patents True green%
A01 Agriculture;forestry; animal husbandry; hunting;trapping; fishing 917,961 246,215 49,224 19.99
A21 Baking; edible doughs 46,654 1,742 383 21.99
A22 Butchering; meat treatment; processing poultry or fish 26,671 1,719 213 12.39
A23 Foods or foodstuffs; treatment thereof, not covered by other classes 340,402 35,332 6,004 16.99
A24 Tobacco; cigars; cigarettes; smokers’ requisites 67,251 3,230 399 12.35
A41 Wearing apparel 150,232 2,533 510 20.13
A42 Headwear 25,377 604 129 21.36
A43 Footwear 74,236 1,211 154 12.72
A44 Haberdashery; jewellery 48,611 863 76 8.81
A45 Hand or travelling articles 172,126 5,334 1,525 28.59
A46 Brushware 32,592 478 75 15.69
A47 Furniture;domestic articles or appliances; coffee mills; spice mills;suction cleaners in general 962,654 34,013 5,857 17.22
A61 Medical or veterinary science; hygiene 1,953,254 99,124 12,838 12.95
A62 Life-saving; fire-fighting 116,182 12,315 2,151 17.47
A63 Sports; games; amusements 376,895 11,390 896 7.87
A99 Subject matter not otherwise provided for in section a 229 33 4 12.12
B01 Physical or chemical processes or apparatus in general 1,332,341 610,168 115,244 18.89
B02 Crushing, pulverising, or disintegrating; preparatory treatment of grain for milling 228,865 36,524 7,200 19.71
B03 Separation of solid materials using liquids or using pneumatic tables or jigs 68,908 29,003 5,716 19.71
B04 Centrifugal apparatus or machines for carrying-out physical or chemical processes 28,460 4,359 590 13.54
B05 Spraying or atomising in general; applying liquids or other fluent materials to surfaces, in general 284,963 31,853 4,729 14.85
B06 Generating or transmitting mechanical vibrations in general 10,698 398 42 10.55
B07 Separating solids from solids; sorting 198,511 18,846 3,011 15.98
B08 Cleaning 437,682 71,352 14,924 20.92
B09 Disposal of solid waste; reclamation of contaminated soil 57,347 55,800 14,642 26.24
B21 Mechanical metal-working without essentially removing material; punching metal 426,356 16,889 1,305 7.73
B22 Casting; powder metallurgy 191,769 23,468 1,587 6.76
B23 Machine tools; metal-working not otherwise provided for 934,873 47,275 3,842 8.13
B24 Grinding; polishing 272,096 14,666 1,182 8.06
B25 Hand tools; portable power-driven tools; tool handles 435,111 15,778 1,193 7.56
B26 Hand cutting tools; cutting; severing 198,219 10,057 1,337 13.29
B27 Working or preserving wood or similar material; nailing or stapling machines in general 83,071 3,920 528 13.47
B28 Working cement, clay, or stone 171,772 12,834 2,064 16.08
B29 Working of plastics; working of substances in a plastic state in general 541,241 52,185 9,481 18.17
B30 Presses 67,789 11,141 2,094 18.8
B31 Making paper articles; working paper 54,608 1,711 249 14.55
B32 Layered products 271,499 29,370 8,669 29.52
B33 Additive manufacturing technology 57,978 14,547 637 4.38
B41 Printing; lining machines; typewriters; stamps 212,671 7,487 1,182 15.79
B42 Bookbinding; albums; files; special printed matter 47,294 870 61 7.01
B43 Writing or drawing implements; bureau accessories 69,636 615 95 15.45
B44 Decorative arts 52,718 1,838 337 18.34
B60 Vehicles in general 952,473 189,175 50,907 26.91
B61 Railways 93,654 9,136 1,133 12.4
B62 Land vehicles for travelling otherwise than on rails 396,455 27,170 4,700 17.3
B63 Ships or other waterborne vessels; related equipment 129,617 32,229 6,311 19.58
  • Note: The table reports patents, green patents, and true green patents distribution across 3-digits CPC classification, together with the % ratio of true green over green - period 2010-2022 (part I).
Table B7: Patents distribution by CPC (part II).
Classification code Description Patents Green patents True green patents True green%
B64 Aircraft; aviation; cosmonautics 151,156 23,886 3,183 13.33
B65 Conveying; packing; storing; handling thin or filamentary material 1,509,769 135,761 19,708 14.52
B66 Hoisting; lifting; hauling 340,274 13,221 1,422 10.76
B67 Opening or closing bottles, jars or similar containers; liquid handling 65,328 2,483 231 9.3
B68 Saddlery; upholstery 4,395 81 9 11.11
B81 Microstructural devices or systems, e.g. micromechanical devices 18,441 945 58 6.14
B82 Nanotechnology 71,511 19,307 3,251 16.84
B99 Subject matter not otherwise provided for in section b 5 1 0 0
C01 Inorganic chemistry 186,032 76,230 16,775 22.01
C02 Treatment of water, waste water, sewage, or sludge 424,551 424,548 154,477 36.39
C03 Glass; mineral or slag wool 84,477 15,830 2,301 14.54
C04 Cements; concrete; artificial stone; ceramics; refractories 123,798 37,915 9,124 24.06
C05 Fertilisers; manufacture thereof 43,951 30,732 7,759 25.25
C06 Explosive; matches 7,239 405 41 10.12
C07 Organic chemistry 393,922 99,795 14,075 14.1
C08 Organic macromolecular compounds; their preparation or chemical working-up 383,504 54,643 11,400 20.86
C09 Dyes; paints, polishes, natural resins, adhesives; compositions not otherwise provided for 288,284 49,263 7,700 15.63
C10 Petroleum, gas or coke industries, technical gases containing carbon monoxide; fuels; lubricant; peat 121,224 50,492 16,968 33.61
C11 Animal or vegetable oils, fats, fatty substances or waxes; fatty acids therefrom; detergents; candles 49,325 8,741 2,350 26.88
C12 Biochemistry; beer; spirits; wine; vinegar; microbiology; enzymology; mutation or genetic engineering 369,038 85,329 17,249 20.21
C13 Sugar industry 4,211 942 498 52.87
C14 Skins; hides; pelts; leather 9,731 422 65 15.4
C21 Metallurgy of iron 120,894 28,226 3,912 13.86
C22 Metallurgy; ferrous or non-ferrous alloys; treatment of alloys or non-ferrous metals 143,994 43,115 7,228 16.76
C23 Coating metallic material; coating material with metallic material; chemical surface treatment 159,252 21,942 3,036 13.84
C25 Electrolytic or electrophoretic processes; apparatus therefor 92,911 25,903 4,850 18.72
C30 Crystal growth 31,655 11,996 1,825 15.21
C40 Combinatorial chemistry; libraries, e.g. chemical libraries, dna libraries 4,867 570 43 7.54
C99 Subject matter not otherwise provided for in section c 6 3 0 0
D01 Natural or man-made threads or fibres; spinning 84,316 6,579 1,104 16.78
D02 Yarns; mechanical finishing of yarns or ropes; warping or beaming 26,954 1,038 134 12.91
D03 Weaving 40,316 1,175 162 13.79
D04 Braiding; lace-making; knitting; trimmings; non-woven fabrics 58,298 3,603 530 14.71
D05 Sewing; embroidering; tufting 42,532 548 30 5.47
D06 Treatment of textiles or the like; laundering; flexible materials not otherwise provided for 209,541 14,122 2,144 15.18
D07 Ropes; cables other than electric 8,943 200 25 12.5
D10 Indexing scheme associated with sublasses of section d, relating to textiles 15,075 1,190 182 15.29
D21 Paper-making; production of cellulose 51,379 7,802 2,098 26.89
D99 Subject matter not otherwise provided for in section d 8 1 1 100
E01 Construction of roads, railways, or bridges 323,633 34,604 6,961 20.12
E02 Hydraulic engineering; foundations; soil shifting 334,875 67,666 10,521 15.55
E03 Water supply; sewerage 193,710 116,767 19,523 16.72
E04 Building 798,420 168,314 65,037 38.64
E05 Locks; keys; window or door fittings; safes 230,626 7,440 1,394 18.74
E06 Doors, windows, shutters, or roller blinds, in general; ladders 205,454 21,416 9,269 43.28
  • Note: The table reports patents, green patents, and true green patents distribution across 3-digits CPC classification, together with the % ratio of true green over green - period 2010-2022 (part II).
Table B8: Patents distribution by CPC (part III).
Classification code Description Patents Green patents True green patents True green%
E21 Earth drilling; mining 343,751 29,455 3,886 13.19
E99 Subject matter not otherwise provided for in section e 10 0 0
F01 Machines or engines in general; engine plants in general; steam engines 194,029 91,170 9,797 10.75
F02 Internal-combustion engines; hot-gas or combustion-product engine plants 222,379 119,569 9,673 8.09
F03 Machines or engines for liquids; wind, spring, weight, or miscellaneous motors 102,508 88,037 18,697 21.24
F04 Engines or pumps 317,489 27,271 3,936 14.43
F05 Indexing schemes relating to engines or pumps in various subclasses of classes f01-f04 61,557 34,853 2,594 7.44
F15 Fluid-pressure actuators; hydraulics or pneumatics in general 86,427 4,544 487 10.72
F16 Engineering elements or units; general measures for effecting and maintaining effective functioning of machines 1,413,612 105,839 14,431 13.63
F17 Storing or distributing gases or liquids 71,948 11,868 1,365 11.5
F21 Lighting 482,334 84,536 29,415 34.8
F22 Steam generation 41,347 14,985 4,945 33
F23 Combustion apparatus; combustion processes 157,204 81,676 18,076 22.13
F24 Heating; ranges; ventilating 551,446 198,789 78,356 39.42
F25 Refrigeration or cooling; combined heating and refrigeration systems; heat pump systems 210,456 48,985 12,638 25.8
F26 Drying 201,684 27,834 6,931 24.9
F27 Furnaces; kilns; ovens; retorts 103,060 32,363 8,476 26.19
F28 Heat exchange in general 143,827 39,344 9,540 24.25
F41 Weapons 52,771 1,087 85 7.82
F42 Ammunition; blasting 34,591 1,260 90 7.14
F99 Subject matter not otherwise provided for in section f 4 0 0
G01 Measuring; testing 2,344,419 171,633 19,988 11.65
G02 Optics 482,028 22,451 2,249 10.02
G03 Photography; cinematography; analogous techniques using waves other than optical waves 246,693 7,362 703 9.55
G04 Horology 34,822 1,531 245 16
G05 Controlling; regulating 367,905 69,208 10,096 14.59
G06 Computing; calculating; counting 1,982,972 187,012 18,944 10.13
G07 Checking-devices 190,290 10,223 2,013 19.69
G08 Signalling 341,630 36,238 6,512 17.97
G09 Educating; cryptography; display; advertising; seals 464,773 27,659 4,892 17.69
G10 Musical instruments; acoustics 120,484 6,012 887 14.75
G11 Information storage 112,154 4,801 153 3.19
G12 Instrument details 4,563 179 20 11.17
G16 Information and communication technology [ict] specially adapted for specific application fields 84,890 8,583 716 8.34
G21 Nuclear physics; nuclear engineering 40,330 22,713 1,258 5.54
G99 Subject matter not otherwise provided for in section g 19 1 0 0
H01 Electric elements 2,001,196 625,938 101,705 16.25
H02 Generation, conversion, or distribution of electric power 1,234,173 436,760 155,010 35.49
H03 Electronic circuitry 175,540 11,015 506 4.59
H04 Electric communication technique 1,725,419 126,469 12,398 9.8
H05 Electric techniques not otherwise provided for 656,099 118,791 15,953 13.43
H10 Semiconductor devices; electric solid-state devices not otherwise provided for 131,510 44,066 6,206 14.08
H99 Subject matter not otherwise provided for in section H 22 3 1 33.33
Y02 Technologies or applications for mitigation or adaptation against climate change 2,439,032 2,439,032 531,075 21.77
Y04 Information or communication technologies having an impact on other technology areas 55,239 55,239 9,858 17.85
Y10 Technical subjects covered by former uspc 172,157 29,600 5,189 17.53
  • Note: The table reports patents, green patents, and true green patents distribution across 3-digits CPC classification, together with the % ratio of true green over green - period 2010-2022 (part III).
Table B9: RCA index by CPC (part I).
Classification code Description RCA green RCA true green
A01 Agriculture;forestry; animal husbandry; hunting;trapping; fishing 1.51 1.43
A21 Baking; edible doughs 0.21 0.22
A22 Butchering; meat treatment; processing poultry or fish 0.36 0.21
A23 Foods or foodstuffs; treatment thereof, not covered by other classes 0.58 0.47
A24 Tobacco; cigars; cigarettes; smokers’ requisites 0.27 0.16
A41 Wearing apparel 0.09 0.09
A42 Headwear 0.13 0.13
A43 Footwear 0.09 0.05
A44 Haberdashery; jewellery 0.1 0.04
A45 Hand or travelling articles 0.17 0.23
A46 Brushware 0.08 0.06
A47 Furniture;domestic articles or appliances; coffee mills; spice mills;suction cleaners in general 0.19 0.16
A61 Medical or veterinary science; hygiene 0.28 0.17
A62 Life-saving; fire-fighting 0.59 0.49
A63 Sports; games; amusements 0.17 0.06
A99 Subject matter not otherwise provided for in section a 0.81 0.46
B01 Physical or chemical processes or apparatus in general 2.64 2.37
B02 Crushing, pulverising, or disintegrating; preparatory treatment of grain for milling 0.89 0.83
B03 Separation of solid materials using liquids or using pneumatic tables or jigs 2.36 2.2
B04 Centrifugal apparatus or machines for carrying-out physical or chemical processes 0.86 0.55
B05 Spraying or atomising in general; applying liquids or other fluent materials to surfaces, in general 0.62 0.44
B06 Generating or transmitting mechanical vibrations in general 0.21 0.1
B07 Separating solids from solids; sorting 0.53 0.4
B08 Cleaning 0.91 0.9
B09 Disposal of solid waste; reclamation of contaminated soil 5.47 6.81
B21 Mechanical metal-working without essentially removing material; punching metal 0.22 0.08
B22 Casting; powder metallurgy 0.68 0.22
B23 Machine tools; metal-working not otherwise provided for 0.28 0.11
B24 Grinding; polishing 0.3 0.11
B25 Hand tools; portable power-driven tools; tool handles 0.2 0.07
B26 Hand cutting tools; cutting; severing 0.28 0.18
B27 Working or preserving wood or similar material; nailing or stapling machines in general 0.26 0.17
B28 Working cement, clay, or stone 0.42 0.32
B29 Working of plastics; working of substances in a plastic state in general 0.54 0.46
B30 Presses 0.92 0.82
B31 Making paper articles; working paper 0.18 0.12
B32 Layered products 0.6 0.85
B33 Additive manufacturing technology 1.4 0.29
B41 Printing; lining machines; typewriters; stamps 0.2 0.15
B42 Bookbinding; albums; files; special printed matter 0.1 0.03
B43 Writing or drawing implements; bureau accessories 0.05 0.04
B44 Decorative arts 0.2 0.17
B60 Vehicles in general 1.11 1.43
B61 Railways 0.55 0.32
B62 Land vehicles for travelling otherwise than on rails 0.38 0.31
B63 Ships or other waterborne vessels; related equipment 1.39 1.29
B64 Aircraft; aviation; cosmonautics 0.88 0.56
  • Note: The table reports RCA index by CPC classification for green patents and true green patents - period 2010-2022 (part I).
Table B10: RCA index by CPC (part II).
Classification code Description RCA green RCA true green
B65 Conveying; packing; storing; handling thin or filamentary material 0.49 0.34
B66 Hoisting; lifting; hauling 0.22 0.11
B67 Opening or closing bottles, jars or similar containers; liquid handling 0.21 0.09
B68 Saddlery; upholstery 0.1 0.05
B81 Microstructural devices or systems, e.g. micromechanical devices 0.29 0.08
B82 Nanotechnology 1.51 1.2
B99 Subject matter not otherwise provided for in section b 1.12 0.0
C01 Inorganic chemistry 2.3 2.4
C02 Treatment of water, waste water, sewage, or sludge 5.77 10.37
C03 Glass; mineral or slag wool 1.05 0.72
C04 Cements; concrete; artificial stone; ceramics; refractories 1.72 1.96
C05 Fertilisers; manufacture thereof 3.92 4.69
C06 Explosive; matches 0.31 0.15
C07 Organic chemistry 1.42 0.95
C08 Organic macromolecular compounds; their preparation or chemical working-up 0.8 0.79
C09 Dyes; paints, polishes, natural resins, adhesives; compositions not otherwise provided for 0.96 0.71
C10 Petroleum, gas or coke industries, technical gases containing carbon monoxide; fuels; lubricant; peat 2.34 3.73
C11 Animal or vegetable oils, fats, fatty substances or waxes; fatty acids therefrom; detergents; candles 0.99 1.26
C12 Biochemistry; beer; spirits; wine; vinegar; microbiology; enzymology; mutation or genetic engineering 1.29 1.24
C13 Sugar industry 1.25 3.13
C14 Skins; hides; pelts; leather 0.24 0.18
C21 Metallurgy of iron 1.31 0.86
C22 Metallurgy; ferrous or non-ferrous alloys; treatment of alloys or non-ferrous metals 1.68 1.33
C23 Coating metallic material; coating material with metallic material; chemical surface treatment 0.77 0.5
C25 Electrolytic or electrophoretic processes; apparatus therefor 1.56 1.38
C30 Crystal growth 2.12 1.53
C40 Combinatorial chemistry; libraries, e.g. chemical libraries, dna libraries 0.66 0.23
C99 Subject matter not otherwise provided for in section c 2.8 0.0
D01 Natural or man-made threads or fibres; spinning 0.44 0.35
D02 Yarns; mechanical finishing of yarns or ropes; warping or beaming 0.22 0.13
D03 Weaving 0.16 0.11
D04 Braiding; lace-making; knitting; trimmings; non-woven fabrics 0.35 0.24
D05 Sewing; embroidering; tufting 0.07 0.02
D06 Treatment of textiles or the like; laundering; flexible materials not otherwise provided for 0.38 0.27
D07 Ropes; cables other than electric 0.13 0.07
D10 Indexing scheme associated with sublasses of section d, relating to textiles 0.44 0.32
D21 Paper-making; production of cellulose 0.85 1.08
D99 Subject matter not otherwise provided for in section d 0.7 3.31
E01 Construction of roads, railways, or bridges 0.6 0.57
E02 Hydraulic engineering; foundations; soil shifting 1.13 0.83
E03 Water supply; sewerage 3.4 2.69
E04 Building 1.18 2.2
E05 Locks; keys; window or door fittings; safes 0.18 0.16
E06 Doors, windows, shutters, or roller blinds, in general; ladders 0.58 1.2
  • Note: The table reports RCA index by CPC classification for green patents and true green patents - period 2010-2022 (part II).
Table B11: RCA index by CPC (part III).
Classification code Description RCA green RCA true green
E21 Earth drilling; mining 0.48 0.3
E99 Subject matter not otherwise provided for in section e 0.0 0.0
F01 Machines or engines in general; engine plants in general; steam engines 2.64 1.34
F02 Internal-combustion engines; hot-gas or combustion-product engine plants 3.03 1.15
F03 Machines or engines for liquids; wind, spring, weight, or miscellaneous motors 4.84 4.87
F04 Engines or pumps 0.48 0.33
F05 Indexing schemes relating to engines or pumps in various subclasses of classes f01-f04 3.18 1.12
F15 Fluid-pressure actuators; hydraulics or pneumatics in general 0.29 0.15
F16 Engineering elements or units; general measures for effecting and maintaining effective functioning of machines 0.41 0.26
F17 Storing or distributing gases or liquids 0.92 0.5
F21 Lighting 0.98 1.62
F22 Steam generation 2.03 3.17
F23 Combustion apparatus; combustion processes 2.92 3.06
F24 Heating; ranges; ventilating 2.03 3.88
F25 Refrigeration or cooling; combined heating and refrigeration systems; heat pump systems 1.3 1.59
F26 Drying 0.77 0.91
F27 Furnaces; kilns; ovens; retorts 1.76 2.18
F28 Heat exchange in general 1.53 1.76
F41 Weapons 0.12 0.04
F42 Ammunition; blasting 0.2 0.07
F99 Subject matter not otherwise provided for in section f 0.0 0.0
G01 Measuring; testing 0.4 0.22
G02 Optics 0.26 0.12
G03 Photography; cinematography; analogous techniques using waves other than optical waves 0.17 0.08
G04 Horology 0.25 0.19
G05 Controlling; regulating 1.05 0.73
G06 Computing; calculating; counting 0.52 0.25
G07 Checking-devices 0.3 0.28
G08 Signalling 0.59 0.5
G09 Educating; cryptography; display; advertising; seals 0.33 0.28
G10 Musical instruments; acoustics 0.28 0.19
G11 Information storage 0.24 0.04
G12 Instrument details 0.22 0.12
G16 Information and communication technology [ict] specially adapted for specific application fields 0.57 0.22
G21 Nuclear physics; nuclear engineering 3.16 0.83
G99 Subject matter not otherwise provided for in section g 0.29 0.0
H01 Electric elements 1.78 1.36
H02 Generation, conversion, or distribution of electric power 2.01 3.52
H03 Electronic circuitry 0.35 0.08
H04 Electric communication technique 0.4 0.19
H05 Electric techniques not otherwise provided for 1.01 0.64
H10 Semiconductor devices; electric solid-state devices not otherwise provided for 1.88 1.25
H99 Subject matter not otherwise provided for in section H 0.76 1.2
Y02 Technologies or applications for mitigation or adaptation against climate change 6.92 7.52
Y04 Information or communication technologies having an impact on other technology areas 5.62 4.75
Y10 Technical subjects covered by former uspc 0.96 0.8
  • Note: The table reports RCA index by CPC classification for green patents and true green patents - period 2010-2022 (part III).